Use references angles to find the exact value of each expression. Sin 300
Answer by mamiya(56)

Use references angles to find the exact value of each expression. Sin 300

sin 300= sin (180+120)
= sin(180)cos(120) + sin(120)cos(180)
= 0*cos(120) + sin(120)(-1)
= -1sin(120)
= -1sin(180-120) since the sinus of two supplementary angles are the same.
so sin 300= -1sin60
= -(sqrt 3)/2